OpenStudy (anonymous):

Part 1: Explain, in complete sentences, which method you would use to solve the following system of equations. (1 point) Part 2: Explain why you chose that method (1 point) Part 3: Provide the solution to the system. (2 points) x – 2y + z = 0 2x – 3y – 4z = –9 x + 2y – 5z = 0

OpenStudy (anonymous):

the method i would use is elimination, so i would try to cancel out y, so i would use both the first equation and the third one then solve to get z=2x-4z=0 then i would plug z back into the second equation so you get 2x-3y-4(1/2x+0=-9 2x-3y-2x=-9 combine like terms you get -3y=-9 divide by three on both sides you get y=3 plug that back into the first equation to get x+z=6 then get x by itself plug that back into the third equation to get z=2 then to get x plug y and z into the second equation you get x=4 Solution you get (4,3,2)

OpenStudy (anonymous):

just want to know if i am right

OpenStudy (ganpat):

looks fine to me.. sorry for late reply..

OpenStudy (anonymous):

its cool and ty

OpenStudy (ganpat):

w/c :)